An electron with total energy E = 0.1 eV is trappped in a finite square well of height 20 eV,

except for the region 0 < x < 2 nm. What is the penetration depth into the classically

forbbidden region x<0?
Answer #1

given

E = 0.1 eV

= 0.1 X 10-19 J

ho = 20 eV

Uo = m g ho

= 9.1 X 10-31 X 9.8 X 20 X 1.6 X 10-19

= 1.426 X 10-48 J

penetration depth = 1/k

k = ( 2 m ( Uo - E ) / \hbar 2 )1/2

\hbar = 2 X 6.67 X 10-34 / 6.28

\hbar = 2.124 X 10-34

k = ( 2 X 9.1 X 10-31 X ( 1.426 X 10-48 - 0.1 X 1.6 X 10-19 ) / ( 2.124 X 10-34 )2 )1/2

k = 803417288.4

penetration depth = 1/k

penetration depth = 1/803417288.4

penetration depth = 1.2446 X 10-9 m

or

penetration depth = 1.2446 nm